# Deciphering Hate: Identifying Hateful Memes and Their Targets

### v1

@@ 5.2 Results
Among the visual approach, the ConVNeXT model obtained the highest score (F1: 0.665). In the case of the textual-only models, Bangla BERT outperformed others (mBERT, XLMR) with an F1 score of 0.644. Notably, this score falls approximately 2.1% short of the best visual model performance, indicating that visual information is more distinguishable in identifying hateful memes. In contrast, the state-of-the-art multimodal model, ALBEF surpassed both unimodal counterparts, attaining the highest F1 score of 0.670. MMBT and CLIP failed to deliver satisfactory results. However, despite ALBEF’s notable performance, our proposed method (DORA) outperforms the best model with an absolute improvement of 4.8% in F1 score.
@@
Table 3 illustrates that in task 2, the textual model (mBERT) achieved the best F1 score of 0.652 among unimodal models. Nonetheless, consistent with our earlier findings, the joint evaluation of multimodal information led to significant enhancements in target identification. In the case of the multimodal models, the ALBEF surpassed the unimodal counterparts, achieving the highest score of 0.668. However, DORA outperforms the best model by 5.2% in terms of F1 score.
@@ 5.3 Ablation Study
For Task 1, it is noteworthy that even in the absence of VF and TF, the model’s performance remains superior (0.691-0.694) compared to the best baseline model, ALBEF (0.670). However, a substantial drop in the F1 score (ranging between 0.654 to 0.675) occurs when the attentive representations are removed, underscoring the significant impact of the dual-co attention mechanism in our proposed approach. Conversely, in the case of target identification, the model exhibits diminished performance when the VF and TF are excluded. Interestingly, removing VGAR and TGAR shows less effect on the performance as the F1 score rises to its highest at 0.686. This implies that, for target identification, VF and TF bear greater significance than attentive representations. However, the integration of all components in DORA results in a notable overall performance boost for both tasks.
@@ 5.4 Transferability and Generalizability of DORA
Table 5 shows the transferability and generalizability of the DORA on two datasets (MUTE Hossain et al. (2022) and EmoffMeme Kumari et al. (2023)) of different languages (i.e., Bengali and Hindi). Both MUTE and EmoffMeme have code-mixed captions in the memes, similar to the BHM dataset. To compare the performance in this experiment, we consider the best baseline model ( ALBEF). Results exhibit that when training and testing are done on the same dataset, DORA exceeds ALBEF by $\approx$ 4-5% in terms of F1 score across all the datasets. This outcome illustrates that DORA can generalize well across languages as well. Similarly, when trained on one dataset and tested on a different one, DORA yields a better score than ALBEF.
One interesting note is that the model trained in the Hindi dataset (EmoffMeme) exhibits poor performance when tested on Bengali datasets. Conversely, the model trained on the Bengali datasets and tested on the Hindi dataset exhibits suboptimal performance (0.64-0.65). This outcome emphasizes the need for a more sophisticated method that can be transferable across various languages. Overall, it can be stated that DORA is generalizable and also transferable across the datasets of the same languages.
@@
Table 5: Transferability of best multimodal baseline and (DORA) on two additional benchmark datasets namely MUTE and EmoffMeme. Here, BEN and HIN indicate the Bengali and Hindi languages respectively. The models are trained on the dataset specified in the rows and tested on the dataset specified in the columns. All the reported scores are weighted F1. The best transferable results are indicated in blue color and the scores in bold denote the best performance when models are trained and tested on the same dataset.
@@ 6 Conclusion
This paper introduced a new large-scale multimodal dataset of 7,148 memes for detecting Bengali hateful memes and their targeted social entities. We also proposed DORA, a multimodal deep neural network for solving the tasks. Experiments on our dataset demonstrate the efficacy of DORA, which outperformed nine state-of-art baselines for two tasks. We further demonstrated the generalizability and transferability of DORA across other datasets of different languages. In the future, we plan to extend the dataset for more domains and languages.
@@
Though the proposed method (DORA) shows superior performance, there still exist some limitations of our work. First, we did not consider the background contexts such as visual entities (i.e., detected objects) and textual entities (i.e., person name, organization name) as external knowledge to the model, which could improve the overall performance. Second, it is likely that in some cases, the DORA may focus on less significant parts of the content while attending to the information. If the dataset contains misleading captions or irrelevant textual information, the attention mechanism might align with those parts of the image that are visually unrelated, producing misleading representations. Incorporating adversarial training could be an interesting future direction to mitigate the generation of such biased multimodal representations. Third, we observed that our method DORA struggled with memes that convey hate implicitly. It appeared to have difficulty correctly interpreting cultural references and context-specific content, leading to additional incorrect predictions. We plan to address this aspect in the future.
@@
Biases: Any biases found in the dataset and model are unintentional. A diverse group of annotators labeled the data following a comprehensive annotation guideline, and all annotations were reviewed to address any potential annotation biases. We randomly collected data from various public social media pages and blogs to reduce any data source biases. Moreover, we used neutral keywords (e.g., Bengali Memes, Bengali Mojar Memes, Funny Memes, Bengali Hashir Memes) not explicitly tied to specific hate themes to mitigate biases toward any specific person, community, or organization. Despite our best efforts, there may be inherent biases in the dataset, a common challenge in the dataset development process.
@@
Reproducibility: We present the details of our experimental setting in Appendix A for the reproducibility of the system. Upon the acceptance of the paper, we will release the source code and the dataset at https://github.com/eftekhar-hossain/Bengali-Hateful-Memes.
@@ Appendix A Experimental Settings
For the experimentation, we use the Google Colab platform. We downloaded the transformer models from the Huggingface [^3] https://huggingface.co/ library and implemented it using the PyTorch Framework. The BNLP [^4] https://github.com/sagorbrur/bnlp and scikit-learn [^5] https://scikit-learn.org/stable/ library has been used for the preprocessing and evaluation measures. We selected the models’ hyperparameter values empirically by examining the performance on the validation set. All the models are compiled using cross_entropy loss function.
@@
On the other hand, in the case of the proposed DORA and its variants, we use the two attention heads in the multi-head co-attention block. During training, the models were optimized using MADGRAD with a learning rate of $2e^{-5}$. We used the batch size of 4 and trained the models for 20 epochs with a learning rate scheduler. We examined the validation set performance to save the best model during training.
@@ Appendix C Additional Data Statistics
Table C.1 presents lexical statistics for the training set meme captions. It is evident that in Task 1, the NHT class exhibits the highest number of unique words(12,428) in comparison to the HT class (8,852). This discrepancy is unsurprising as the NHT class has the largest number of instances among all the classes. In the case of Task 2, the TI class got the highest count of unique words (7,150) while the TS class features the lowest (727). The average caption length remains consistent at 13 words for most classes, with exceptions in TC (14) and TS (12). Figure C.1 displays a histogram illustrating caption length across different classes. The distribution reveals that the majority of captions fall within the 5 to 30-word range. Only the TI and NHT classes contain captions with lengths exceeding 30 words. We further analyzed the captions to quantify word overlap across different classes. Specifically, we computed the Jaccard similarity (JS) Jain et al. (2017) score between the top 400 most common words of different classes. The Jaccard score between each pair of classes is presented in Table C.2. We observed a substantial JS score of 0.51 between the HT and NHT classes, indicating a significant overlap in the words of these two classes. Regarding the target categories, the TI and TC pair exhibited the highest JS score (0.34), while the scores for the other categories remained below 0.20.
@@
Task 1 HT 28477
8852
13.45
NHT 50344
12428
13.82
Task 2 TI 21583
7150
13.29
TO 2159
1362
13.49
TC 3694
2017
14.83
TS 1041
727
12.24
Table C.1: Lexical analysis of captions in the training set in terms of total words, total unique words, and average caption length across categories.
@@
The presence/absence of facial images is an important component in any meme. Therefore, we analyze the faces present in a meme. To perform this analysis, we employed the deepFaceSerengil and Ozpinar (2021)library. It allowed us to determine whether a given meme contains any faces. If a face is detected, we also extract information regarding the associated gender and age. Our findings reveal that approximately 34.12% of the memes do not contain any faces. Among this group, 33.28% belong to the ’Hate’ class. Out of the total memes, 53.29% feature male faces, with 35.3% from the ’Hate’ class and 64.64% from the ’Not-Hate’ class. Meanwhile, 12.53% feature female faces, with 50.83% within the ’Hate’ class and 49.17% within the ’Not-Hate’ class. On average, the detected ages of both males and females in the memes hover around 31 years. A key observation is the prevalence of male faces in the ’Hate’ class, indicating that males are common targets of hateful content in the Bengali community.

### v2

@@ 5.2 Results
Among the visual approach, the ConVNeXT model obtained the highest score (F1: 0.665). In the case of the textual-only models, Bangla BERT outperformed others (mBERT, XLMR) with an F1 score of 0.644. Notably, this score falls approximately 2.1% short of the best visual model performance, indicating that visual information is more distinguishable in identifying hateful memes. In contrast, ALBEF surpassed both unimodal counterparts in the state-of-the-art multimodal model, attaining the highest F1 score of 0.670. MMBT and CLIP failed to deliver satisfactory results. However, despite ALBEF’s notable performance, our proposed method (DORA) outperforms the best model with an absolute improvement of 4.8% in F1 score.
@@
Table 3 illustrates that task 2’s textual model (mBERT) achieved the best F1 score of 0.652 among unimodal models. Nonetheless, consistent with our earlier findings, the joint evaluation of multimodal information led to significant enhancements in target identification. In the case of the multimodal models, the ALBEF surpassed the unimodal counterparts, achieving the highest score of 0.668. However, DORA outperforms the best model by 5.2% in terms of F1 score.
@@ 5.3 Ablation Study
For Task 1, it is noteworthy that even in the absence of VF and TF, the model’s performance remains superior (0.691-0.694) compared to the best baseline model, ALBEF (0.670). However, a substantial drop in the F1 score (ranging between 0.654 to 0.675) occurs when the attentive representations are removed, underscoring the significant impact of the dual-co attention mechanism in our proposed approach. Conversely, in the case of target identification, the model exhibits diminished performance when the VF and TF are excluded. Interestingly, removing VGAR and TGAR shows less effect on the performance as the F1 score rises to its highest at 0.686. This implies that, for target identification, VF and TF bear greater significance than attentive representations. However, integrating all components in DORA results in a notable overall performance boost for both tasks.
@@ 5.4 Transferability and Generalizability of DORA
Table 5 shows the transferability and generalizability of the DORA on two datasets (MUTE Hossain et al. (2022) and EmoffMeme Kumari et al. (2023)) of different languages (i.e., Bengali and Hindi). Both MUTE and EmoffMeme have code-mixed captions in the memes, similar to the BHM dataset. To compare the performance in this experiment, we consider the best baseline model ( ALBEF). Results exhibit that when training and testing are done on the same dataset, DORA exceeds ALBEF by $\approx$ 4-5% in terms of F1 score across all the datasets. This outcome illustrates that DORA can also generalize well across languages. Similarly, when trained on one dataset and tested on a different one, DORA yields a better score than ALBEF.
Interestingly, the model trained in the Hindi dataset (EmoffMeme) exhibits poor performance when tested on Bengali datasets. Conversely, the model trained on the Bengali datasets and tested on the Hindi dataset exhibits suboptimal performance (0.64-0.65). This outcome emphasizes the need for a more sophisticated method that can be transferable across various languages. Overall, it can be stated that DORA is generalizable and also transferable across the datasets of the same languages.
@@
Table 5: Transferability of best multimodal baseline and (DORA) on two additional benchmark datasets namely MUTE and EmoffMeme. Here, BEN and HIN indicate the Bengali and Hindi languages, respectively. The models are trained on the dataset specified in the rows and tested on the dataset specified in the columns. All the reported scores are weighted F1. The best transferable results are indicated in blue, and the scores in bold denote the best performance when models are trained and tested on the same dataset.
@@ 6 Conclusion
This paper introduced a new large-scale multimodal dataset of 7,148 memes for detecting Bengali hateful memes and their targeted social entities. We also proposed DORA, a multimodal deep neural network for solving the tasks. Experiments on our dataset demonstrate the efficacy of DORA, which outperformed nine state-of-art baselines for two tasks. We further demonstrated the generalizability and transferability of DORA across other datasets of different languages. We plan to extend the dataset for more domains and languages in the future.
@@
Though the proposed method (DORA) shows superior performance, there still exist some limitations of our work. First, we did not consider the background contexts, such as visual entities (i.e., detected objects) and textual entities (i.e., person name, organization name), as external knowledge to the model, which could improve the overall performance. Second, it is likely that in some cases, the DORA may focus on less significant parts of the content while attending to the information. If the dataset contains misleading captions or irrelevant textual information, the attention mechanism might align with those parts of the image that are visually unrelated, producing misleading representations. Incorporating adversarial training could be an interesting future direction to mitigate the generation of such biased multimodal representations. Third, we observed that our method DORA struggled with memes that convey hate implicitly. It appeared to have difficulty correctly interpreting cultural references and context-specific content, leading to additional incorrect predictions. We plan to address this aspect in the future.
@@
Biases: Any biases found in the dataset and model are unintentional. A diverse group of annotators labeled the data following a comprehensive annotation guideline, and all annotations were reviewed to address any potential annotation biases. We randomly collected data from various public social media pages and blogs to reduce data source biases. Moreover, we used neutral keywords (e.g., Bengali Memes, Bengali Mojar Memes, Funny Memes, Bengali Hashir Memes) not explicitly tied to specific hate themes to mitigate biases toward any specific person, community, or organization. Despite our best efforts, there may be inherent biases in the dataset, a common challenge in the dataset development process.
@@
Reproducibility: We present the details of our experimental setting in Appendix A for the system’s reproducibility. We will release the source code and the dataset at https://github.com/eftekhar-hossain/Bengali-Hateful-Memes upon accepting the paper.
@@ Appendix A Experimental Settings
For the experiment, we used the Google Colab platform. We downloaded the transformer models from the Huggingface [^3] https://huggingface.co/ library and implemented it using the PyTorch Framework. The BNLP [^4] https://github.com/sagorbrur/bnlp and scikit-learn [^5] https://scikit-learn.org/stable/ library has been used for the preprocessing and evaluation measures. We empirically selected the models’ hyperparameter values by examining the validation set’s performance. All the models are compiled using cross_entropy loss function.
@@
On the other hand, in the case of the proposed DORA and its variants, we use the two attention heads in the multi-head co-attention block. During training, the models were optimized using MADGRAD with a $2e^{-5}$ learning rate. We used the batch size of 4 and trained the models for 20 epochs with a learning rate scheduler. We examined the validation set performance to save the best model during training.
@@ Appendix C Additional Data Statistics
Table C.1 presents lexical statistics for the training set meme captions. In Task 1, the NHT class exhibits the highest number of unique words(12,428) compared to the HT class (8,852). This discrepancy is unsurprising as the NHT class has the largest number of instances among all the classes. In the case of Task 2, the TI class got the highest count of unique words (7,150), while the TS class features the lowest (727). The average caption length remains consistent at 13 words for most classes, with exceptions in TC (14) and TS (12). Figure C.1 displays a histogram illustrating caption length across different classes. The distribution reveals that most captions fall within the 5 to 30-word range. Only the TI and NHT classes contain captions with lengths exceeding 30 words. We further analyzed the captions to quantify word overlap across different classes. Specifically, we computed the Jaccard similarity (JS) Jain et al. (2017) score between the top 400 most common words of different classes. The Jaccard score between each pair of classes is presented in Table C.2. We observed a substantial JS score of 0.51 between the HT and NHT classes, indicating a significant overlap in the words of these two classes. Regarding the target categories, the TI and TC pair exhibited the highest JS score (0.34), while the scores for the other categories remained below 0.20.
@@
Task 1 HT 28477 8852 13.45
NHT 50344 12428 13.82
Task 2 TI 21583 7150 13.29
TO 2159 1362 13.49
TC 3694 2017 14.83
TS 1041 727 12.24
Table C.1: Lexical analysis of captions in training set in terms of total words, total unique words, and average caption length across categories.
@@
The presence/absence of facial images is an important component in any meme. Therefore, we analyze the faces present in a meme. We employed the deepFaceSerengil and Ozpinar (2021)library to perform this analysis. It allowed us to determine whether a given meme contains any faces. If a face is detected, we also extract information regarding the associated gender and age. Our findings reveal that approximately 34.12% of the memes contain no faces. Among this group, 33.28% belong to the ’Hate’ class. Of the total memes, 53.29% feature male faces, with 35.3% from the ’Hate’ class and 64.64% from the ’Not-Hate’ class. Meanwhile, 12.53% feature female faces, with 50.83% within the ’Hate’ class and 49.17% within the ’Not-Hate’ class. On average, the detected ages of males and females in the memes hover around 31 years. A key observation is the prevalence of male faces in the ’Hate’ class, indicating that males are common targets of hateful content in the Bengali community.